# Preference-Conditioned Gradient Variations for Multi-Objective Quality-Diversity

### v1

@@ IV MOME-P2C
Figure 3: Overview of mome-p2c algorithm. Pareto Fronts are stored in each cell of a map-elites grid. At each iteration, a batch of solutions are selected, undergo variation and are added back to the grid based on their performance and crowding-distances. As solutions are evaluated, environment transitions are gathered in a replay buffer and used to train preference-conditioned networks. These networks are used with a preference sampler to perform preference-conditioned pg updates.
@@
In this section, we introduce Multi-Objective Map-Elites with Preference-Conditioned Policy-Gradient and Crowding Mechanisms (mome-p2c), a new moqd algorithm that learns a single, preference-conditioned actor-critic framework to provide policy-gradient variations in tasks that can be framed as MDP. The algorithm inherits the core framework of existing moqd methods, which involves maintaining a Pareto front within each feature cell of a map-elites grid and follows a map-elites loop of selection, variation, and addition for a given budget. Building on the approach of mome-pgx, our method not only employs traditional genetic variation operators but also integrates policy gradient mutations that improve sample-efficiency, particularly in high-dimensional search spaces. Similar to mome-pgx, mome-p2c adopts crowding-based selection, which strategically directs exploration towards less explored areas of the search space and also utilizes crowding-based addition mechanisms to promote a continuous distribution of solutions along the Pareto front. Distinct from mome-pgx, which operates with a separate actor-critic framework for each objective function, mome-p2c innovates by employing a singular, preference-conditioned actor-critic. This design streamlines preference-conditioned policy gradient variation updates to genotypes, significantly reducing the memory requirements of the algorithm and making it more scalable to problems with a higher number of objectives. Furthermore, mome-p2c leverages the preference-conditioned actor by injecting it into the main population. A visual representation of the algorithm is depicted in Figure 3, and the accompanying pseudo-code is provided in Algorithm 1. Detailed descriptions of each component of mome-p2c are available in the following sections.
@@
  • mome archive $\mathcal{A}$ and total number of iterations $N$
@@
[…]
@@
     $(f_{1},...,f_{m},d,\text{transitions})\leftarrow\text{evaluate}(\pi_{\tilde{\theta_{1}}},...,\pi_{\tilde{\theta_{b+b_{a}}}})$
@@
     $\mathcal{A}\leftarrow\text{crowding\_addition}(\tilde{\theta_{b+1}},...,\tilde{\theta_{b+b_{a}}})$
@@
IV-A Crowding-based Selection and Addition
In mome-p2c, following mome-pgx [25], we choose to use biased selection and addition mechanisms. In particular, when selecting parent solutions from the grid, we first select a cell with uniform probability and then select an individual from the cell’s Pareto front with probability proportional to its crowding distance. As defined in nsga-ii [19], the crowding distance of a solution is defined as the average Manhattan distance between itself and its k-nearest neighbours, in objective space. In mome-pgx, it was shown that biasing solutions in this manner provides an effective method for guiding the optimisation process toward under-explored regions of the solution space.
Similarly, we also use a crowding-informed addition mechanisms to replace solutions on the Pareto front. It is important to note that all moqd methods we consider use a fixed maximum size for the Pareto front of each cell. This is done in order to exploit the parallelism capabilities of recent hardware advances [48, 49] and consequently affords many thousands of evaluations in a short period of time. However, if a solution is added to a Pareto front that is at already maximum capacity, another solution must also necessarily be removed. In mome-p2c, following from mome-pgx, we remove the solution with the minimum crowding distance in order to sparsity of solutions on the front.
Further details regarding the crowding-based mechanisms can be found in the mome-pgx paper [25]. To justify these selection and addition mechanisms are still valid for mome-p2c, we include an ablation of these crowding mechanisms in our ablation study (see Section VI-B).
IV-B Preference-Conditioned Actor-Critic
In mome-pgx, a separate actor-critic framework was used to find a policy $\pi$ that marginally maximised the expected sum of rewards $\bm{J}^{i}(\pi)=\mathbb{E}_{\pi}[\sum_{t}\bm{r_{t}}^{i}]$ for each objective $i=1,...,m$. However, in mome-p2c, we do not require a separate actor-critic framework for each objective function. Instead, we use a single actor-critic framework that aims to find a single actor policy to maximise $J(\pi,\bm{\omega})=\mathbb{E}_{\pi}[\sum_{t}\bm{\omega}^{\intercal}\bm{r}_{t}]$ for any given preference $\bm{\omega}$.
@@ IV-B Preference-Conditioned Actor-Critic
$\begin{split}Q^{\pi}(s,a|\bm{\omega})&=\mathbb{E}_{\pi(\cdot|\bm{\omega})}\left[\sum_{t=0}^{T}\gamma^{t}\bm{\omega}^{\intercal}\bm{r}(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]\\ &=\bm{\omega}^{\intercal}\mathbb{E}_{\pi(\cdot|\bm{\omega})}\left[\sum_{t=0}^{T}\gamma^{t}\bm{r}(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]\\ &=\bm{\omega}^{\intercal}\bm{Q}^{\pi}(s,a)\end{split}$ (8)
Here, $\bm{Q}^{\pi}(s,a|\bm{\omega})$ denotes the preference-conditioned vectorised action-value function. Equation 8 demonstrates that that the we can estimate the preference-conditioned action-value function by training a critic $\bm{Q}_{\psi}(s,a|\bm{\omega})\to\mathbb{R}^{m}$ to predict the vectorised action-value function and then weighting its output by the preference. To train this critic network, we modify the target td3 algorithm given in Equation 5 to be:
$y=\bm{\omega}^{\intercal}\bm{r}(s_{t},a_{t})+\gamma\min_{i=1,2}\bm{\omega}^{\intercal}\bm{Q}_{\psi_{i}}(s_{t+1},\pi_{\phi^{\prime}}(s_{t+1}|\bm{\omega})+\epsilon|\bm{\omega})$ (9)
@@
$\begin{split}\nabla_{\phi}J(\phi,\bm{\omega})&=\bm{\omega}^{\intercal}\mathbb{E}\big{[}\,\nabla_{\phi}\,\pi_{\phi}(s|\bm{\omega})\,\nabla_{a}\bm{Q_{\psi}}(s,a|\bm{\omega})\,|\,_{a=\pi_{\phi}(s|\bm{\omega})}\big{]}\\ &=\mathbb{E}\big{[}\,\nabla_{\phi}\,\pi_{\phi}(s|\bm{\omega})\,\nabla_{a}\bm{\omega}^{\intercal}\bm{Q_{\psi}}(s,a|\bm{\omega})\,|\,_{a=\pi_{\phi}(s|\bm{\omega})}\big{]}\end{split}$ (10)
The updates of the actor and critic networks, given by Equation 9 and Equation 10, depend on the value of the preference $\bm{\omega}$. In mome-p2c, for each sampled transition, we uniformly sample a preference and use this to form a single policy gradient update. Since the preference vector assumes that each of the objectives are scaled equally, we normalise the reward values using a running mean and variance throughout the algorithm. Solutions are stored and added to the archive based on unnormalised fitnesses.
IV-C Preference-Conditioned Policy Gradient Variation
Given the preference-conditioned actor-critic framework described in Section IV-B, we can form preference-conditioned pg variations on solutions in the archive. In mome-p2c at each iteration, we select $b_{p}$ solutions from the archive and perform n of preference-conditioned policy gradient steps via:
$\nabla_{\theta}J(\theta,\bm{\omega})=\mathbb{E}\big{[}\,\nabla_{\theta}\,\pi_{\theta}(s)\,\nabla_{a}\mathbf{Q_{\psi_{1}}}(s,a|\bm{\omega})\,|\,_{a=\pi_{\theta}(s)}\big{]}$ (11)
The pg update given by Equation 11 depends on a preference vector $\bm{\omega}$. However, it is not straightforward to determine the best strategy for choosing the value of this vector. In this work, we use the term “pg preference sampler” to refer to present the strategy we use for determining the preference that the pg variation is conditioned on (illustrated in Figure 3). In mome-p2c, we choose the pg preference sampler to simply be a random uniform sampler as we found this to be a simple, yet effective strategy. We examine other choices for the pg preference sampler in our ablation study (Section VI-B).
IV-D Actor Injection
In pga-me, mome-pgx and other gradient-based qd methods, the actor policy has the same shape as the policies stored in the map-elites grid and so can be regularly injected into the main offspring batch as a genotype, with no additional cost to the main algorithm. However, in mome-p2c, the policies in the map-elites grid only take the current state $s_{t}$ as input, whereas the preference-conditioned actor takes the state concatenated with a preference $[s_{t},\bm{\omega}]$ as input. Therefore, the actor has a different architecture to the policies so cannot be added to the repertoire. In this work, we take a similar approach to the one taken by dcrl-me to inject the conditioned actor within the population [47].
@@ IV-D Actor Injection
$W\begin{bmatrix}s_{t}\\ \bm{\omega}\end{bmatrix}+B$ $=\begin{bmatrix}W_{s}&W_{\bm{\omega}}\end{bmatrix}\begin{bmatrix}s_{t}\\ \bm{\omega}\end{bmatrix}+B$ (12)
@@
V Experimental Setup
In this section, we describe the evaluation tasks, baselines and metrics we use to evaluate our approach. Importantly, we introduce two new tri-objective moqd tasks, Ant-3 and Hopper-3, which allow us to evaluate the capabilities of different moqd approaches to scale to a larger number of objectives.. We also introduce two new moqd metrics, moqd-sparsity-score and global-sparsity, which we argue are important ways to assess whethermoqd algorithms are able to achieve smooth sets of trade-offs. These new tasks and metrics form two key contributions of this work.
TABLE I: Summary of evaluation tasks.
@@ V-A Evaluation Tasks
We evaluate our approach on six continuous control robotics tasks, which are summarised in Table I. In these tasks, solutions correspond to the parameters of closed-loop neural network controllers which determine the torque commands given to each of the robot’s joints. We use four robot morphologies from the Brax suite [50]. In all of the tasks, the feature is the proportion of time that each of the robot’s legs spends in contact with the ground. Using this characterisation, solutions that have diverse features will exhibit different gaits [8, 21].
@@
In four of the tasks (Ant-2, HalfCheetah-2, Hopper-2, Walker-2) the aim is to maximise the forward velocity of the robot while minimising its energy consumption [21, 25]. However, we also introduce two tri-objective moqd environments: Ant-3 and Hopper-3. In the Ant-3 task, the objectives are the robot’s $x$-velocity, $y$-velocity and energy consumption. Hence the goal is to discover controllers that lead to different gaits, and for each of these gaits to find controllers that travel in different directions while minimising the energy cost. In the Hopper-3 task, the rewards correspond to the robot’s forward velocity, torso height and energy consumption. The corresponding moqd goal is to therefore find solutions which have different gaits, and for each of these gaits to find controllers that make the hopper jump to different heights or travel forward, while minimising the energy cost. We designed these tasks as they present interesting and realistic objectives, and also provide opportunity to compare moqd algorithms on tasks with $m>2$.
V-B Baselines
@@
We compare mome-p2c to five baselines: mome-pgx, mome, pga-me, nsga-ii and spea2. mome-pgx and mome are both moqd algorithms so are straightforward to evaluate. In pga-me, we convert the multiple objectives into a single one by adding them. To ensure all algorithms have equal population sizes, if we use a grid of $k$ cells with maximum Pareto length of $|\mathcal{P}|$ for moqd methods, we use $k\times|\mathcal{P}|$ cells for pga-me and a population size of $k\times|\mathcal{P}|$ for nsga-ii and spea2. To report metrics for pga-me, nsga-ii and spea2, we use a passive archive with the same structure as the moqd methods. At each iteration, we fill the passive archive with solutions found by the algorithm and then calculate metrics on these archives. Importantly, the passive archives do not interact within the primary algorithmic loop, ensuring that there is no effect on the behaviour of the baseline algorithms.
V-C Metrics

### v2

@@ 4. MOME-P2C
Figure 3. Overview of mome-p2c algorithm. Non-dominated Fronts are stored in each cell of a map-elites grid. At each iteration, a batch of solutions are selected, undergo variation and are added back to the grid based on their performance and crowding-distances. As solutions are evaluated, environment transitions are gathered in a replay buffer and used to train preference-conditioned networks. These networks are used with a preference sampler to perform preference-conditioned pg updates.
@@
In this section, we introduce Multi-Objective Map-Elites with Preference-Conditioned Policy-Gradient and Crowding Mechanisms (mome-p2c), a new moqd algorithm that learns a single, preference-conditioned actor-critic framework to provide policy-gradient variations in tasks that can be framed as MDP. The algorithm inherits the core framework of existing moqd methods, which involves maintaining a non-dominated front within each feature cell of a map-elites grid and follows a map-elites loop of selection, variation, and addition for a given budget. Building on the approach of mome-pgx, our method not only employs traditional genetic variation operators but also integrates policy gradient mutations that improve sample-efficiency, particularly in high-dimensional search spaces. Similar to mome-pgx, mome-p2c adopts crowding-based selection, which strategically directs exploration towards less explored areas of the search space and also utilizes crowding-based addition mechanisms to promote a continuous distribution of solutions along the non-dominated front. Distinct from mome-pgx, which operates with a separate actor-critic framework for each objective function, mome-p2c innovates by employing a singular, preference-conditioned actor-critic. This design streamlines preference-conditioned policy gradient variation updates to genotypes, reducing the memory requirements of the algorithm and making it more scalable to problems with a higher number of objectives (a numerical comparison is shown in Appendix C). Furthermore, mome-p2c leverages the preference-conditioned actor by injecting it into the main population. A visual representation of the algorithm is depicted in Figure 3, and the accompanying pseudo-code is provided in Algorithm 1. Detailed descriptions of each component of mome-p2c are available in the following sections.
@@
  • mome archive $\mathcal{A}$ and total number of iterations NN
@@
  $\bm{\omega}_{b_{p}+1},...,\bm{\omega}_{b_{p}+b_{a}}\leftarrow\text{actor\_sampler}$
@@
[…]
@@
[…]
@@
4.1. Crowding-based Selection and Addition
In mome-p2c, following mome-pgx (Janmohamed et al., 2023), we choose to use biased selection and addition mechanisms. In particular, when selecting parent solutions from the grid, we first select a cell with uniform probability and then select an individual from the cell’s non-dominated front with probability proportional to its crowding distance. As defined in nsga-ii (Deb et al., 2000), the crowding distance of a solution is defined as the average Manhattan distance between itself and its nearest neighbours on either side of itself in objective space. In mome-pgx, it was shown that biasing solutions in this manner provides an effective method for guiding the optimisation process toward under-explored regions of the solution space.
Similarly, we also use a crowding-informed addition mechanisms to replace solutions on the non-dominated front. It is important to note that all moqd methods we consider use a fixed maximum size for the non-dominated front of each cell. This is done in order to exploit the parallelism capabilities of recent hardware advances (Lim et al., 2022; Chalumeau et al., 2024) and consequently affords many thousands of evaluations in a short period of time. However, if a solution is added to a non-dominated front that is at already maximum capacity, another solution must also necessarily be removed. In mome-p2c, following from mome-pgx, we remove the solution with the minimum crowding distance in order to increase the sparsity of solutions on the front.
Further details regarding the crowding-based mechanisms can be found in the mome-pgx paper (Janmohamed et al., 2023). While we acknowledge that crowding distance may lose effectiveness in many-objective settings (Zheng et al., 2024; Köppen and Yoshida, 2007), we adopted it here as a simple and efficient choice consistent with mome-pgx, and note that it still yields strong results (we include an ablation of the crowding mechanisms in our ablation study Section 6.2.2). Nevertheless, we note that exploring alternative approaches for preserving diversity of the non-dominated front with a higher number of objectives remains an interesting avenue for future work.
4.2. Preference-Conditioned Actor-Critic
In mome-pgx, a separate actor-critic framework was used to find a policy $\pi$ that marginally maximised the expected sum of rewards $\bm{J}^{i}(\pi)=\mathbb{E}_{\pi}[\sum_{t}\bm{r_{t}}^{i}]$ for each objective i=1,,mi=1,...,m. However, in mome-p2c, we do not require a separate actor-critic framework for each objective function. Instead, we use a single actor-critic framework that aims to find a single actor policy to maximise $J(\pi,\bm{\omega})=\mathbb{E}_{\pi}[\sum_{t}\bm{\omega}^{\intercal}\bm{r}_{t}]$ for any given preference $\bm{\omega}$.
@@ 4.2. Preference-Conditioned Actor-Critic
(8) $\begin{split}Q^{\pi}(s,a|\bm{\omega})&=\mathbb{E}_{\pi(\cdot|\bm{\omega})}\left[\sum_{t=0}^{T}\gamma^{t}\bm{\omega}^{\intercal}\bm{r}(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]\\ &=\bm{\omega}^{\intercal}\mathbb{E}_{\pi(\cdot|\bm{\omega})}\left[\sum_{t=0}^{T}\gamma^{t}\bm{r}(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]\\ &=\bm{\omega}^{\intercal}\bm{Q}^{\pi}(s,a)\end{split}$
Here, $\bm{Q}^{\pi}(s,a|\bm{\omega})$ denotes the preference-conditioned vectorised action-value function. Equation 8 demonstrates that the we can estimate the preference-conditioned action-value function by training a critic $\bm{Q}_{\psi}(s,a|\bm{\omega})\to\mathbb{R}^{m}$ to predict the vectorised action-value function and then weighting its output by the preference. To train this critic network, we modify the target td3 algorithm given in Equation 5 to be:
(9) $y=\bm{\omega}^{\intercal}\bm{r}(s_{t},a_{t})+\gamma\min_{i=1,2}\bm{\omega}^{\intercal}\bm{Q}_{\psi_{i}}(s_{t+1},\pi_{\phi^{\prime}}(s_{t+1}|\bm{\omega})+\epsilon)$
@@
(10) $\begin{split}\nabla_{\phi}J(\phi,\bm{\omega})&=\bm{\omega}^{\intercal}\mathbb{E}\big[\,\nabla_{\phi}\,\pi_{\phi}(s|\bm{\omega})\,\nabla_{a}\bm{Q_{\psi}}(s,a|\bm{\omega})\,|\,_{a=\pi_{\phi}(s|\bm{\omega})}\big]\\ &=\mathbb{E}\big[\,\nabla_{\phi}\,\pi_{\phi}(s|\bm{\omega})\,\nabla_{a}\bm{\omega}^{\intercal}\bm{Q_{\psi}}(s,a|\bm{\omega})\,|\,_{a=\pi_{\phi}(s|\bm{\omega})}\big]\end{split}$
The updates of the actor and critic networks, given by Equation 9 and Equation 10, depend on the value of the preference $\bm{\omega}$. In mome-p2c, for each sampled transition, we uniformly sample a preference and use this to form a single policy gradient update. Since the preference vector assumes that each of the objectives are scaled equally, we normalise the reward values using a running mean and variance throughout the algorithm. Solutions are stored and added to the archive based on unnormalised fitnesses for interpretability and comparison to other methods.
4.3. Preference-Conditioned Policy Gradient Variation
Given the preference-conditioned actor-critic framework described in Section 4.2, we can form preference-conditioned pg variations on solutions in the archive. In mome-p2c at each iteration, we select $b_{p}$ solutions from the archive and perform n preference-conditioned policy gradient steps. For a given solution $\theta$ which parametrises a policy, $\pi_{\theta(s)}$ we update the parameters via:
(11) $\nabla_{\theta}J(\theta,\bm{\omega})=\mathbb{E}\big[\,\nabla_{\theta}\,\pi_{\theta}(s)\,\nabla_{a}\mathbf{Q_{\psi_{1}}}(s,a|\bm{\omega})\,|\,_{a=\pi_{\theta}(s)}\big]$
The pg update given by Equation 11 depends on a preference vector $\bm{\omega}$. However, it is not straightforward to determine the best strategy for choosing the value of this vector. In this work, we use the term “pg preference sampler” to refer to the strategy we use for determining the preference that the pg variation is conditioned on (illustrated in Figure 3). In mome-p2c, we choose the pg preference sampler to simply be a random uniform sampler as we found this to be a simple, yet effective strategy. We examine other choices for the pg preference sampler in our ablation study (Section 6.2).
4.4. Actor Injection
In pga-me, mome-pgx and other gradient-based qd methods, the actor policy has the same shape as the policies stored in the map-elites grid and so can be regularly injected into the main offspring batch as a genotype, with no additional cost to the main algorithm. However, in mome-p2c, the policies in the map-elites grid only take the current state $s_{t}$ as input, whereas the preference-conditioned actor takes the state concatenated with a preference $s_{t},\bm{\omega}]$ as input. Therefore, the actor has a different architecture to the policies so cannot be added to the repertoire. In this work, we take a similar approach to the one taken by dcrl-me to inject the conditioned actor within the population (Faldor et al., 2024).
@@ 4.4. Actor Injection
(12) $W\begin{bmatrix}s_{t}\\ \bm{\omega}\end{bmatrix}+B$ $=\begin{bmatrix}W_{s}&W_{\bm{\omega}}\end{bmatrix}\begin{bmatrix}s_{t}\\ \bm{\omega}\end{bmatrix}+B$
@@
5. Experimental Setup
In this section, we describe the evaluation tasks, baselines and metrics we use to evaluate our approach. We introduce two new tri-objective moqd tasks, ant-3 and hopper-3, which allow us to evaluate the capabilities of different moqd approaches to scale to a larger number of objectives. We also introduce two new moqd metrics (moqd-sparsity-score and global-sparsity) which we argue are important ways to assess whether moqd algorithms are able to achieve smooth sets of trade-offs. These new tasks and metrics form two key contributions of this work.
Table 1. Summary of evaluation tasks.
@@ 5.1. Evaluation Tasks
We evaluate our approach on six continuous control robotics tasks, which are summarised in Table 1. In these tasks, solutions correspond to the parameters of closed-loop neural network controllers which determine the torque commands given to each of the robot’s joints. In all tasks, the neural network controllers consist of two hidden layers of 64 neurons. The exact number of parameters depends on the observation and action dimensionalities of each environment, as summarised in Table 1. We use four robot morphologies from the Brax suite (Freeman et al., 2021). In all of the tasks, the feature is the proportion of time that each of the robot’s legs spends in contact with the ground. Given a robot with $L$ legs (summarised in Table 1), the $L$-dimensional feature space is calculated as:
(13) $\phi_{i}(x)=\frac{1}{T}\sum_{t=1}^{T}c_{i,t}\,\,\,\text{for}\,\,i=1,...,L\,.$
Here, $c_{i,t}$ is $1$ when the leg is in contact with the ground at time step $t$, otherwise it is 0. Using this characterisation, solutions that have diverse features will exhibit different gaits (Cully et al., 2015; Pierrot et al., 2022b).
@@
In four of the tasks (ant-2, halfcheetah-2, hopper-2, walker-2) the aim is to maximise the forward velocity of the robot while minimising its energy consumption (Pierrot et al., 2022b; Janmohamed et al., 2023). However, we also introduce two tri-objective moqd environments: ant-3 and hopper-3. In the ant-3 task, the objectives are the robot’s $x$-velocity, $y$-velocity and energy consumption. Hence the goal is to discover controllers that lead to different gaits, and for each of these gaits to find controllers that travel in different directions while minimising the energy cost. In the hopper-3 task, the rewards correspond to the robot’s forward velocity, torso height and energy consumption. The corresponding moqd goal is to therefore find solutions which have different gaits, and for each of these gaits to find controllers that make the hopper jump to different heights or travel forward, while minimising the energy cost. These tasks were inspired by those in the pg-morl paper (Xu et al., 2020) but extended to be moqd problems. We believe that they present interesting and realistic objectives, and also provide opportunity to compare moqd algorithms on tasks with $m>2$.
5.2. Baselines
We compare mome-p2c to seven baselines: mome-pgx, mome, pga-me, nsga-ii and spea2. Despite our best efforts, we were unable to implement pg-morl as a morl baseline, as the available code had compatibility issues with the QDax framework (Chalumeau et al., 2024) that we use for parallel evaluations of policies via Brax (Freeman et al., 2021), making pg-morl more than 250 times slower than the other baselines.
mome-pgx and mome are both moqd algorithms so are straightforward to evaluate. However, nsga-ii and spea2 are population-based multi-objective Evolutionary Algorithms (see Section 3.1) which do not explicitly seek diversity and pga-me is a mono-objective qd algorithm (see Section 3.3). To evaluate pga-me we convert the multiple objectives into a single one by adding them. To ensure that pga-me had the same archive size as moqd methods, if we use a grid of $k$ cells with maximum non-dominated front length of $|\mathcal{P}|$ for moqd methods, we use $k\times|\mathcal{P}|$ cells for pga-me. To ensure that our comparisons are not confounded by differences in archive size, we also include an additional pga-me baseline configured with the same number of grid cells as the MOQD methods (see Appendix D). By contrast, using a large population size for nsga-ii and spea2 can reduce selection pressure and limit the number of generations, potentially hindering performance. To address this, we run two versions of both baselines: (1) the nsga-ii and spea2 baselines, which use population sizes as close to $k\times|\mathcal{P}|$ that were feasible within our GPU compute limits, and (2) the nsga-ii-small and spea2-small baselines, which use the same batch size as the qd baselines but run for more generations to match the total number of evaluations.
@@ 5.2. Baselines
To report metrics for pga-me, nsga-ii, nsga-ii-small, spea2 and spea2-small we use a passive archive with the same structure as the moqd methods. At each iteration, we fill the passive archive with solutions found by the algorithm and then calculate metrics on these archives. Importantly, the passive archives do not interact within the primary algorithmic loop, ensuring that there is no effect on the behaviour of the baseline algorithms.
5.3. Metrics